# Physics-Informed Weakly Supervised Learning for Interatomic Potentials

### v1

@@ B.2 Baseline Methods
Fig. 1 (a) illustrates the corresponding approach [17], which computes the energy of a perturbed atomic configuration using a Taylor expansion based on reference energy and atomic force labels. This approach was originally applied to train MLIPs without explicit force labels.
B.3 Description of the Data Sets
ANI-1x Data Set. The ANI-1x data set is a heterogeneous molecular data set and includes 63,865 organic molecules (with chemical elements H, C, N, and O) whose size ranges from 4 to 64 atoms [19]. The ML model requires learning total energies and atomic forces for various molecules and their conformations. Total energies and atomic forces are obtained through DFT calculations.
TiO2 Data Set. Titanium dioxide (TiO2) is an industrially relevant and well-studied material. TiO2 dataset includes 7815 bulk structures of several TiO2 phases whose reference energies and forces are obtained through DFT calculations [26]. The number of atoms in a single configuration ranges from 6 to 95.
rMD17 Data Set. The rMD17 data set includes ten small organic molecules, including 100,000 configurations obtained by running MD simulations for each [51]. The ML model requires learning the total energies and atomic forces for each molecule. In this revised version of the MD17 data set, the molecules are taken from the original MD17 data set [46, 50]. However, the energies and forces are recalculated at the PBE/def2-SVP level of theory using very tight SCF convergence and a very dense DFT integration grid.
LMNTO Data Set. The Li-Mo-Ni-Ti oxide (LMNTO) is of technological significance as a potential high-capacity positive electrode material for lithium-ion batteries. It exhibits substitutional disorder, with Li, Mo, Ni, and Ti all sharing the same sublattice. This data set includes LMNTO with the composition Li8Mo2Ni7Ti7O32 and configurations obtained from an MD simulation, resulting in approximately 2600 structures in total [17].
Appendix C Differences in Gradients for Physics-Informed Losses
@@
In contrast, for the PITC loss in Eq. (3) we obtain
@@
Appendix D Experiments
D.1 Benchmark Results
@@
Additional Results for ANI-1x. Table A4 provides results for ANI-1x data set and a training set sizes of 50 or 10,000; see also Fig. 2. The table demonstrates a considerable reduction of energy and force RMSEs for models trained using small training data set sizes of 50 configurations. Furthermore, we find around 5 to 25 % error reduction for a larger training set size of 10,000, indicating the effectiveness of the PIWSL method for relatively large training set sizes.
@@
Additional Results for SchNet Applied to TiO2. In Table 2, we set the mini-batch size to 32 for training the SchNet model. This adjustment was made because training SchNet with a small mini-batch size of four increases RMSE values with a growing training data set size. Table A5 demonstrates the performance of the SchNet model for a mini-batch size of four. Table A6 provides the results obtained for the SchNet model with a mini-batch size of four for the following training set sizes: 100, 200, 500, and 1000. This figure demonstrates that SchNet, with a mini-batch size of four, reaches its best performed with $N_{\rm train}=200$. These results indicate the difficulty of learning training data statistics from small mini-batch information, probably due to the limited expressive power of SchNet.
@@
Results for LMNTO. Table A7 presents RMSE errors for LMNTO [17]. PIWSL shows the error reduction for most cases for this benchmark data set, especially for small training set sizes (i.e., a training set of 100 configurations).
@@
Table A8: Energy (E) and force (F) root-mean-square errors (RMSEs) for the rMD17 data set. We have chosen benzene, naphthalene, and aspirin for our experiments. The results are obtained by averaging over three independent runs. Energy RMSE is given in kcal/mol, while force RMSE is in kcal/mol/Å.
$N_{\mathrm{train}}=100$ $N_{\mathrm{train}}=1000$
Dataset Model Baseline NoisyNode PIWSL Baseline NoisyNode PIWSL
Schnet E 0.23 $\pm$ 0.00 0.58 $\pm$0.00 0.23 $\pm$0.00 0.17 $\pm$ 0.00 0.32 $\pm$ 0.00 0.17 $\pm$ 0.00
F 2.32 $\pm$ 0.00 3.61 $\pm$ 0.00 2.32 $\pm$ 0.00 1.27 $\pm$ 0.00 2.51 $\pm$ 0.00 1.27 $\pm$ 0.00
PaiNN E 0.90 $\pm$ 0.02 2.29 $\pm$ 0.50 0.89 $\pm$0.03 0.47 $\pm$ 0.03 0.75 $\pm$ 0.02 0.49 $\pm$ 0.03
F 0.57 $\pm$ 0.00 5.33 $\pm$ 0.16 0.57 $\pm$ 0.00 0.23 $\pm$ 0.00 2.50 $\pm$ 0.00 0.30 $\pm$ 0.00
Benzene SpinConv E 2.27 $\pm$ 0.09 2.32 $\pm$ 0.00 1.61 $\pm$0.28 0.90 $\pm$ 0.12 2.35 $\pm$ 0.01 1.07 $\pm$ 0.00
F 0.61 $\pm$ 0.01 3.56 $\pm$ 0.00 0.65 $\pm$ 0.01 0.39 $\pm$ 0.00 2.33 $\pm$ 0.00 0.43 $\pm$ 0.00
($n_{\rm atom}=12$) eSCN E 0.59 $\pm$ 0.01 3.47 $\pm$ 0.04 0.58 $\pm$ 0.03 0.20 $\pm$ 0.00 1.01 $\pm$ 0.01 0.19 $\pm$ 0.00
F 0.74 $\pm$ 0.01 8.43 $\pm$ 0.18 0.75 $\pm$ 0.02 0.14 $\pm$ 0.00 2.99 $\pm$ 0.01 0.14 $\pm$ 0.00
Equiformer E 1.55 $\pm$ 0.01 2.08 $\pm$ 0.01 1.52 $\pm$ 0.01 0.281 $\pm$ 0.01 1.68 $\pm$ 0.02 0.276 $\pm$ 0.01
F 0.72 $\pm$ 0.00 10.32 $\pm$ 0.04 0.72 $\pm$ 0.01 0.15 $\pm$ 0.00 2.89 $\pm$ 0.00 0.13 $\pm$ 0.00
Schnet E 1.41 $\pm$ 0.00 1.92 $\pm$ 0.00 1.41 $\pm$ 0.00 1.05 $\pm$ 0.00 1.49 $\pm$ 0.00 1.05 $\pm$ 0.00
F 5.76 $\pm$ 0.00 5.96 $\pm$ 0.00 5.76 $\pm$ 0.00 3.80 $\pm$ 0.00 4.08 $\pm$ 0.00 3.80 $\pm$ 0.00
PaiNN E 3.63 $\pm$ 0.01 5.13 $\pm$ 0.06 3.54 $\pm$ 0.02 1.37 $\pm$ 0.02 2.22 $\pm$ 0.04 1.33 $\pm$ 0.01
F 1.98 $\pm$ 0.01 10.99 $\pm$ 0.05 1.99 $\pm$ 0.00 0.72 $\pm$ 0.00 2.56 $\pm$ 0.01 0.72 $\pm$ 0.00
Naphthalene SpinConv E 2.96 $\pm$ 0.22 5.73 $\pm$ 0.00 2.88 $\pm$ 0.02 1.80 $\pm$ 0.02 3.39 $\pm$ 0.00 2.40 $\pm$ 0.18
F 2.04 $\pm$ 0.01 3.91 $\pm$ 0.00 1.99 $\pm$ 0.01 0.97 $\pm$ 0.00 2.46 $\pm$ 0.00 0.96 $\pm$ 0.00
($n_{\rm atom}=18$) eSCN E 2.07 $\pm$ 0.03 7.63 $\pm$ 0.05 2.12 $\pm$ 0.01 0.56 $\pm$ 0.01 2.15 $\pm$ 0.29 0.58 $\pm$ 0.01
F 2.28 $\pm$ 0.01 9.68 $\pm$ 0.23 2.32 $\pm$ 0.23 0.42 $\pm$ 0.01 2.86 $\pm$ 0.03 0.42 $\pm$ 0.01
Equiformer E 4.37 $\pm$ 0.03 5.70 $\pm$ 0.05 4.27 $\pm$ 0.01 0.71 $\pm$ 0.02 3.70 $\pm$ 0.10 0.72 $\pm$ 0.02
F 1.93 $\pm$ 0.03 12.73 $\pm$ 0.06 1.89 $\pm$ 0.00 0.43 $\pm$ 0.02 3.20 $\pm$ 0.02 0.38 $\pm$ 0.06
Schnet E 3.76 $\pm$ 0.00 3.56 $\pm$ 0.00 3.74 $\pm$ 0.00 2.77 $\pm$ 0.00 3.08 $\pm$ 0.00 2.77$\pm$ 0.00
F 12.32 $\pm$ 0.00 11.59 $\pm$ 0.00 12.20 $\pm$ 0.00 6.63 $\pm$ 0.00 7.03 $\pm$ 0.00 6.63 $\pm$ 0.00
PaiNN E 6.55 $\pm$ 0.03 9.36 $\pm$ 0.08 5.64 $\pm$ 0.02 4.07 $\pm$ 0.01 4.10 $\pm$ 0.01 3.99$\pm$ 0.01
F 7.38 $\pm$ 0.02 20.37 $\pm$ 0.04 7.36 $\pm$ 0.03 2.17 $\pm$ 0.00 2.17 $\pm$ 0.00 2.16 $\pm$ 0.01
Aspirin SpinConv E 5.71 $\pm$ 0.04 6.11 $\pm$ 0.00 5.03 $\pm$ 0.01 4.04 $\pm$ 0.09 4.12 $\pm$ 0.06 3.42 $\pm$ 0.20
F 8.68 $\pm$ 0.03 10.17 $\pm$ 0.00 8.94 $\pm$ 0.02 1.88 $\pm$ 0.00 1.89 $\pm$ 0.00 1.83$\pm$ 0.01
($n_{\rm atom}=21$) eSCN E 5.14 $\pm$ 0.02 6.44 $\pm$ 0.04 4.82$\pm$ 0.13 1.28 $\pm$ 0.03 1.28 $\pm$ 0.02 1.29 $\pm$ 0.01
F 6.14 $\pm$ 0.03 13.88$\pm$ 0.16 6.10 $\pm$ 0.03 1.30 $\pm$ 0.01 1.29$\pm$ 0.02 1.30 $\pm$ 0.01
Equiformer E 4.79 $\pm$ 0.02 5.75 $\pm$ 0.08 4.66 $\pm$ 0.04 1.83 $\pm$ 0.04 1.83 $\pm$ 0.06 1.75 $\pm$ 0.01
F 4.86 $\pm$ 0.03 16.80 $\pm$ 0.05 4.86 $\pm$ 0.03 1.00 $\pm$ 0.03 1.00 $\pm$ 0.00 0.94 $\pm$ 0.08
Molecular Dynamics Trajectories (rMD17). Table A8 analyzes the effect of the PIWSL when training using conformations of a single small molecule, different from the heterogeneous ANI-1x data set. For this purpose, we have chosen the benzene ($n_{\rm atom}=12$), naphthalene ($n_{\rm atom}=18$), and aspirin ($n_{\rm atom}=21$) molecules because these represent molecules of different sizes. The results in Table A8 indicate that our approach is still effective in this scenario. However, the PIWSL performance in the case of benzene has nearly no gain. This observation can be attributed to only a small variation of atomic coordinates in the benzene data set, simplifying the learning task. We also note that the obtained results for SchNet and PaiNN are somewhat worse than originally reported [28, 34], attributed to the modified implementation in the Open Catalyst project code and the use of the force-branch instead of the gradient-based forces. The effect of the gradient-based forces is investigated in Section D.9.
D.2 Training Time Analysis
@@
Table A9 provides training times measured for experiments with and without PIWSL. The training time is measured for a single training epoch and is averaged over five epochs in total. The experiments were performed using 1000 training configurations from the ANI-1x data set. We used a mini-batch of six. The table indicates that PIWSL increases the training time by around two to three compared to the baseline (due to the additional gradient calculations). We emphasize that the PIWSL approach only alters training time; the inference time is unaffected.
D.3 Different Configurations for the Physics-Informed Spatial-Consistency Loss
Table A10: Results for different configurations of the PISC loss. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set [19]. Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations. The case 1, 2, and 3 correspond to Eq. (4), Eq. (A8) and Eq. (A9), respectively.
@@ D.3 Different Configurations for the Physics-Informed Spatial-Consistency Loss
where the point at $\mathbf{r}+\delta\mathbf{r}$ is the point where PIRC loss is imposed (see Eq. (3)). The results are provided in Table A10 and indicate that Eq. (4) (Case 1) shows a better performance than the other cases for both the potential energy and the force predictions. In this study, we used the ANI-1x data set with 1000 training samples different from the one used to train the model used in the main body to avoid overfitting on the test data set. For the coefficient of the PITC and PISC losses, we used $0.1$ and $0.001$ with $\epsilon_{\rm max}=0.01$.
D.4 C–H Potential Energy Profile of Aspirin

### v2

@@ C.2 Baseline Methods
Figure 1 (a) illustrates the corresponding approach (Cooper et al., 2020), which computes the energy of a perturbed atomic configuration using a Taylor expansion based on reference energy and atomic force labels. This approach was originally applied to train MLIPs without explicit force labels.
C.3 Description of the Data Sets
ANI-1x Data Set. The ANI-1x data set is a heterogeneous molecular data set and includes 63,865 organic molecules (with chemical elements H, C, N, and O) whose size ranges from 4 to 64 atoms (Smith et al., 2020). The ML model requires learning total energies and atomic forces for various molecules and their conformations. Total energies and atomic forces are obtained through DFT calculations.
TiO2 Data Set. Titanium dioxide (TiO2) is an industrially relevant and well-studied material. TiO2 dataset includes 7815 bulk structures of several TiO2 phases whose reference energies and forces are obtained through DFT calculations (Artrith & Urban, 2016). The number of atoms in a single configuration ranges from 6 to 95.
rMD17 Data Set. The rMD17 data set includes ten small organic molecules, including 100,000 configurations obtained by running MD simulations for each (Christensen & von Lilienfeld, 2020). The ML model requires learning the total energies and atomic forces for each molecule. In this revised version of the MD17 data set, the molecules are taken from the original MD17 data set (Chmiela et al., 2017, 2018). However, the energies and forces are recalculated at the PBE/def2-SVP level of theory using very tight SCF convergence and a very dense DFT integration grid.
LMNTO Data Set. The Li-Mo-Ni-Ti oxide (LMNTO) is of technological significance as a potential high-capacity positive electrode material for lithium-ion batteries. It exhibits substitutional disorder, with Li, Mo, Ni, and Ti all sharing the same sublattice. This data set includes LMNTO with the composition Li8Mo2Ni7Ti7O32 and configurations obtained from an MD simulation, resulting in approximately 2600 structures in total (Cooper et al., 2020).
MD22 Data Set. The MD22 data set  (Chmiela et al., 2023) includes seven larger organic molecules, such as a small peptide and a double-walled nanotube, whose size ranges from 42 to 370 atoms. The data set consists of MD trajectories sampled at temperatures between 400 and 500 K. The ML model requires learning the total energies and atomic forces for each molecule. The energies and forces are calculated at the PBE+MBD level of theory.
Appendix D Differences in Gradients for Physics-Informed Losses
@@
In contrast, for the PITC loss in Eq. (4.1) we obtain
@@
Appendix E Experiments
E.1 Benchmark Results
@@
Table A5: Energy and force erorrs of PaiNN model trained on the ANI-1x data set with 100,000, 1,000,000, and 5,000,000 samples. The results are obtained by averaging over three independent runs. Energy errors are given in kcal/mol, while force errors are in kcal/mol/Å.
$N_{\rm train}$ Force MAE Force RMSE Energy MAE Energy RMSE
Baseline 100,000 0.92 $\pm$ 0.00 3.70 $\pm$ 0.01 4.28 $\pm$ 0.15 6.14 $\pm$ 0.21
PIWSL 0.91 $\pm$ 0.01 3.72 $\pm$ 0.04 4.07 $\pm$ 0.17 5.83 $\pm$ 0.20
Baseline 1,000,000 0.67 $\pm$ 0.00 2.74 $\pm$ 0.01 4.90 $\pm$ 0.23 6.56 $\pm$ 0.26
PIWSL 0.68 $\pm$ 0.00 2.77 $\pm$ 0.04 4.48 $\pm$ 0.05 6.06 $\pm$ 0.01
Baseline 5,000,000a 0.53 2.18 3.94 5.24
PIWSL 0.55 2.24 3.25 4.75
a Because of the computational cost, we performed only one training in the case of 5,000,000 training samples.
Additional Results for ANI-1x. Table A4 provides results for ANI-1x data set and a training set sizes of 50 or 10,000; see also figure 2. The table demonstrates a considerable reduction of energy and force RMSEs for models trained using small training data set sizes of 50 configurations. Furthermore, we find around 5 to 25 % error reduction for a larger training set size of 10,000, indicating the effectiveness of the PIWSL method for relatively large training set sizes. Finally, we provide a result of PaiNN model trained on ANI-1x data set with 100,000, 1,000,000, and 5,000,000 samples in Table A5, which demonstrates that PIWSL still improves the performance around 5% to 10%of the energy RMSE.
@@
Additional Results for SchNet Applied to TiO2. In Table 2, we set the mini-batch size to 32 for training the SchNet model. This adjustment was made because training SchNet with a small mini-batch size of four increases RMSE values with a growing training data set size. Table A6 demonstrates the performance of the SchNet model for a mini-batch size of four. Table A7 provides the results obtained for the SchNet model with a mini-batch size of four for the following training set sizes: 100, 200, 500, and 1000. This figure demonstrates that SchNet, with a mini-batch size of four, reaches its best performed with $N_{\rm train}=200$. These results indicate the difficulty of learning training data statistics from small mini-batches, probably due to the limited expressive power of SchNet.
@@
Results for LMNTO. Table A8 presents RMSE errors for LMNTO (Cooper et al., 2020). PIWSL shows the error reduction for most cases for this benchmark data set, especially for small training set sizes (i.e., a training set of 100 configurations).
@@
Table A9: Energy (E) and force (F) mean-absolute errors (MAEs) for the MD22 data set. We have chosen buckyball catcher for our experiments. The results are obtained by averaging over three independent runs. Energy MAE is given in kcal/mol/atom, while force MAE is in kcal/mol/Å.
$N_{\mathrm{train}}=50$ $N_{\mathrm{train}}=600$
Dataset Model Baseline PIWSL Baseline PIWSL
Buckyball catcher MACE E 1.046 $\pm$ 0.095 0.745 $\pm$ 0.031 0.587 $\pm$ 0.107 0.548 $\pm$ 0.015
F 0.294 $\pm$ 0.002 0.290 $\pm$ 0.002 0.082 $\pm$ 0.001 0.082 $\pm$ 0.001
Molecular Dynamics Trajectories for Large Molecules – MD22. Table A9 evaluates the impact of PIWSL on data sets containing conformations of a single large molecule. For this purpose, we selected the buckyball catcher molecule with $n_{\rm atom}=148$ atoms. To demonstrate the applicability of PIWSL, we trained a MACE model (Batatia et al., 2022) to prove that PIWSL enhances even the performance of a recent state-of-the-art model. The model structure and training configuration were based on those provided in the official repository [^6] https://mace-docs.readthedocs.io/en/latest/examples/training_examples.html with slight modifications: “max_$L$” was set to one and mini-batch size was adjusted to four. Following the original training setup (Chmiela et al., 2023), we randomly sampled 600 configurations for the training dataset and 400 for the validation dataset and retained the remaining 5102 configurations for testing. To further validate PIWSL’s effectiveness in sparse data scenarios, we prepared a smaller training dataset comprising only 50 configurations while keeping the validation dataset size unchanged. The model was trained for 450 and 800 epochs for the 600-sample and 50-sample training datasets, respectively. The results presented in Table A9 demonstrate that our approach remains effective on average, particularly in the sparse data regime. In this study, the coefficient of the PITC and PISC losses are set as 0.01 and 0.001 with $\epsilon_{\rm max}=0.01$.
In addition to the experiment of training from scratch, we conducted an additional study using the MACE foundation model to evaluate whether PIWSL could be effectively applied in the context of foundation models with finetuning. Specifically, we utilized two foundation models: MACE-MP ("large") (Batatia et al., 2023), a universal model, and MACE-OFF ("large_off") (Kovács et al., 2023), a model designed for organic force fields. Training was performed on the buckyball catcher molecule from the MD22 dataset. To simulate finetuning on a smaller dataset, the models were trained using 50 samples, consistent with the previous experiment. The official settings from the respective repositories were used, with adjustments made to the mini-batch size (set to 4) and the number of epochs (set to 100).
The results, presented in Table 4, demonstrate that PIWSL is effective for the fine-tuning of foundation models. Moreover, the result indicates that the MACE-MP model provides worse results than MACE-OFF, emphasizing the importance of selecting an appropriate pre-trained model. For this experiment, we employed the second-order PITC loss and the two-point consistency loss defined in Eq. (A14) and Eq. (A11), with the coefficients set to 0.08.
E.2 Training Time Analysis
@@
Table A11: Training time comparison for experiments with and without PIWSL on MACE-OFF. We measure the time required for a single training epoch and provide the results obtained as an average over five epochs. We use 950 configurations from the MD17-CC data set. All training times are provided in seconds.
Baseline PIWSL
MACE-OFF 18.5 31.7
Table A10 provides training times measured for experiments with and without PIWSL. The training time is measured for a single training epoch and is averaged over five epochs in total. The experiments were performed using 1000 training configurations from the ANI-1x data set. We used a mini-batch of six. The table indicates that PIWSL increases the training time by a factor of two to three compared to the baseline (due to the additional gradient calculations). This is primarily because our PITC and PISC losses effectively double or triple the number of data labels, resulting in a proportional increase in training time due to the expanded set of training labels. This issue can be mitigated by employing the 2pt-PISC loss introduced in section E.3.2, which eliminates the need to estimate the third conformation. Table A11 presents the average training times per epoch for MACE-OFF over ten epochs, as discussed in subsection 5.4. These results indicate that even the MACE-OFF (large) model requires less than twice the training time while achieving approximately a 40% reduction in error. We emphasize that the PIWSL approach only alters training time; the inference time is unaffected.
E.3 Different Configurations for the Physics-Informed Spatial-Consistency Loss
E.3.1 Triangle-Based
Table A12: Results for different configurations of the PISC loss. The presented numerical values are the root mean square errors (RMSEs) for the ANI-1x data set (Smith et al., 2020). Energy (in kcal/mol) and force (in kcal/mol/Å) errors are obtained by averaging over three independent runs. All models are trained using 1000 configurations. The case 1, 2, and 3 correspond to Eq. (4), Eq. (A8) and Eq. (A9), respectively.
@@ E.3.1 Triangle-Based
where the point at $\mathbf{r}+\delta\mathbf{r}$ is the point where PIRC loss is imposed (see Eq. (4.1)). The results are provided in Table A12 and indicate that Eq. (4) (Case 1) shows a better performance than the other cases for both the potential energy and the force predictions. In this study, we used the ANI-1x data set with 1000 training samples different from the one used to train the model used in the main body to avoid overfitting on the test data set. For the coefficient of the PITC and PISC losses, we used $0.1$ and $0.001$ with $\epsilon_{\rm max}=0.01$.
E.3.2 Further Variations
Two-Point Spacial Consistency.
The flexibility of the PISC loss allows us to explore additional forms of spatial consistency. For example, instead of using a triangular configuration, we can impose spatial consistency between two points at $\mathbf{r}$ and $\mathbf{r}+\delta\mathbf{r}$, leading to the following expression:
$\begin{split}L_{\mathrm{PISC,2pt}}\left(\mathcal{S};\bm{\theta}\right)&=\ell\left(E\left(\mathcal{S};\bm{\theta}\right),E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\langle-\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)\rangle\right).\end{split}$ (A10)
While not thoroughly investigated, we empirically observed that this loss delivers competitive performance when applied with the same coefficient value as the PITC loss.
Two-Point Spacial Consistency with Label.
Since the training sample includes a label, we can utilize this label information instead of predicting the potential energy of the original conformation to enforce two-point spatial consistency.
$\begin{split}L_{\mathrm{PISC,2ptwl}}\left(\mathcal{S};\bm{\theta}\right)&=\ell\left(E_{\mathcal{S}}^{\rm ref},E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\langle-\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)\rangle\right).\end{split}$ (A11)
We observed that this loss function performs well when applying the MACE model to datasets without force labels, as discussed in subsection 5.4.
Second-order Term Consideration for PITC Loss.
In Eq. (4.1), we considered only the first-order Taylor expansion. Here, we introduce a straightforward approach to approximately account for the second-order term. Using the energy and forces at ${\bf r}+\delta\mathbf{r}$, the second-order derivative of the potential energy can be approximated using the explicit finite difference method as:
$\delta\mathbf{r}^{\mu}\delta\mathbf{r}^{\nu}\partial_{\mu}\partial_{\nu}E=\delta\mathbf{r}^{\mu}\delta\mathbf{r}^{\nu}\partial_{\mu}F_{\nu}\simeq\langle\delta\mathbf{r},\mathbf{F}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)-\mathbf{F}\left(\mathcal{S};\bm{\theta}\right)\rangle,$ (A12)
Then, Eq. (4.1) with the above second-order term can be written as:
$E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)$ $\approx E\left(\mathcal{S};\bm{\theta}\right)$
$-\sum_{i=1}^{N_{\mathrm{at}}}\left[\langle\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)\rangle+k_{\rm 2nd}\langle\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)-\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)\rangle\right]+\mathcal{O}\left(\|\delta\mathbf{r}\|^{3}\right),$
$=E\left(\mathcal{S};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\langle\delta\mathbf{r}_{i},(1-k_{\rm 2nd})\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)+k_{\rm 2nd}\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)\rangle+\mathcal{O}\left(\|\delta\mathbf{r}\|^{3}\right),$ (A13)
where $k_{\rm 2nd}$ is a numeric parameter that controls the contribution of the second-order term. Note that setting $k_{\rm 2nd}=0.5$ recovers the exact second-order expression, while $k_{\rm 2nd}=0$ reverts to the first-order approximation.
Using Eq. (E.3.2), the expression of the second-order PITC loss is given as:
$\begin{split}L_{\mathrm{PITC}}\left(\mathcal{S};\bm{\theta}\right)&=\ell\Big{(}E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right),E\left(\mathcal{S};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\langle\delta\mathbf{r}_{i},(1-k_{\rm 2nd})\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)+k_{\rm 2nd}\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)\rangle\Big{)},\end{split}$ (A14)
Note that the second-order term becomes significant under the following conditions: (1) the accuracy of the MLIP model surpasses the contribution of the first-order term, and (2) the original conformation is near the equilibrium state. The latter can be understood as follows. Considering a spring model as a two-body interaction, the potential energy can be expressed as:
$V({\bf r})=\frac{k_{\rm bond}}{2}({\bf r}-{\bf r}_{0})^{2},$ (A15)
where $k_{\rm bond}$ is the constant characterizing the strength of the two-body interaction, and ${\rm r_{0}}$ denotes the equilibrium bond length. Introducing a small perturbation $\delta\mathbf{r}$ in the bond length, the potential energy becomes:
$V({\bf r}+\delta\mathbf{r})=\frac{k_{\rm bond}}{2}({\bf r}-{\bf r}_{0}+\delta\mathbf{r})^{2}=V({\bf r})+2k_{\rm bond}\langle{\bf r}-{\bf r}_{0},\delta\mathbf{r}\rangle+\frac{k_{\rm bond}}{2}\delta\mathbf{r}^{2}.$ (A16)
At the equilibrium state (${\bf r}={\bf r}_{0}$), the above equation demonstrates that the second-order term becomes dominant. Consequently, the second-order accuracy of the PITC loss function becomes crucial in such scenarios.
PITS for Curl of Forces
[…]
E.4 Detailed Setups for Qualitative Analysis
E.4.1 C–H Potential Energy Profile of Aspirin